OpenStudy (anonymous):

Can someone check my answer i got C.) Solve for x in the following equation 5x^2-56=0 A.) x= -7 ; 8/5 B.)x=-7 ;8/5 C.) x=7 ;-8/5 D.)x=7;8/5

OpenStudy (anonymous):

There's something wrong. 5x^2 = 56 so x^2 = 56/5 = 11.2, so x = sqrt11.2 = 3.347. What do your answers ABCD mean? What does -7; 8/5 mean?